OpenStudy (anonymous):

Determine whether the given ordered pair is a solution of the system of equation (5, -10) 3x=25-y 6x=60+3y

OpenStudy (anonymous):

plug in the 5 and -10. 3(5)=25-(-10)......15=35..NO 6(5)=60+3(-10).....30=30YES